# MoPINNEnKF: Iterative Model Inference using generic-PINN-based ensemble Kalman filter

### v1

@@ 3.2.3 Inverse Problem
Table 8 shows the estimated $\alpha$ values of theTFMDWEs and their $L^{1}$ error in the inverse problem setting for (1). ADAM-PINN, (2). NSGA-III-PINN, and (3). MoPINNEnKF with different noise level of observational data. It is shown that in all cases, MoPINNEnKF is more accurate in estimating the missing parameters in TFMDWEs than ADAM-PINN and NSGA-III-PINN. Similar as in the forward problem setting, the NSGA-III-PINN shows more accurate estimations than ADAM-PINN, as the NSGA-III algorithm more effectively balances multiple objectives within the loss function, thereby partially mitigating the adverse effects of noisy observations even in the missing parameters. This is also reflected in the figures 1012 as well as Figure 13 which shows at different time instances (top: $t=0.5$, bottom: $t=1$), the true solution (black) and inverse problem solutions obtained using ADAM-PINN (red), NSGA-III-PINN (blue), and MoPINNEnKF (green) for the time-fractional mixed diffusion-wave equations (TFMDWEs) with different noise levels. The results show that MoPINNEnKF (green) outperforms both ADAM-PINN (red) and NSGA-III-PINN (blue) in terms of estimating parameters and PDE solutions.
Figure 10: Comparison of (a) the true solution and (b) the MoPINNEnKF solution as well as the $L_{1}$ errors of (c) the ADAM-PINN and (d) the MoPINNEnKF for the time-fractional mixed diffusion-wave equations (TFMDWEs) in inverse problem. Observation noise level is $20\%$.
Figure 11: Comparison of (a) the true solution and (b) the MoPINNEnKF solution as well as the $L_{1}$ errors of (c) the ADAM-PINN and (d) the MoPINNEnKF for the time-fractional mixed diffusion-wave equations (TFMDWEs) in inverse problem. Observation noise level is $50\%$.
Figure 12: Comparison of (a) the true solution and (b) the MoPINNEnKF solution as well as the $L_{1}$ errors of (c) the ADAM-PINN and (d) the MoPINNEnKF for the time-fractional mixed diffusion-wave equations (TFMDWEs) in inverse problem. Observation noise level is $80\%$.
@@
Figure 13: Comparison at different time instances (top: $t=0.5$, bottom: $t=0.7$) between the true solution (black) and inverse problem solutions obtained using ADAM-PINN (red), NSGA-III-PINN (blue), and MoPINNEnKF (green) for the time-fractional mixed diffusion-wave equations (TFMDWEs) under varying noise levels: Panel (a) — $20\%$ noise, Panel (b) — $50\%$ noise, and Panel (c) — $80\%$ noise.
@@
MoPINNEnKF 0.5713 0.0713
Table 8: Comparison of PINN with Adam optimizer (ADAM-PINN), PINN with NSGA-III optimizer (NSGA-III-PINN) and the proposed integrated method (MoPINNEnKF) for estimating the fractional term $\alpha$ in the TFMDWEs. For reference, the true value of the fractional term is $\alpha=0.5$.
@@ 4 Conclusion and Future Work
In this paper, we introduce a novel MoPINNEnKF framework that integrates the physics informed neural network (PINN) with ensemble Kalman filter (EnKF) with the NSGA-III multi-objective optimizer. This framework can be used to address both forward problem and inverse problem with only limited noisy data in the context of partial differential equations (PDE). The MoPINNEnKF framework has several advantages: (i) The framework utilizes the multi-objective optimizer NSGA-III to find an optimal cluster of PINNs, and (ii) The uncertainty within this cluster, inherent from the NSGA-III optimization, can be integrated with observational noisy data using the ensemble Kalman filter. This process can then be iteratively applied to update the data loss during PINN training. (iii) The framework can be applied in both the forward and inverse problems in solving PDEs.
Building on the MoPINNEnKF framework’s effectiveness in solving forward and inverse problems with noisy data, several future directions necessitate further investigation. First, in the inverse problem setting, the unknown parameters are treated as trainable variables within the neural networks and are implicitly embedded in the PDE residual loss of the PINN. To quantify the uncertainty of the inferred parameters, these parameters together with the independent variables can be used as inputs to the neural network, allowing it to be trained across a wide range of physical parameters [23, 20]. Second, a Bayesian PINN framework for both inverse and forward models, as proposed in [60], can be integrated with the current framework. This integration could potentially improve the performance of the model in the presence of highly noisy data, as the Bayesian approach naturally quantifies the uncertainties arising from scattered noisy data. Third, the MoPINNEnKF framework can be extended to a continual learning framework, allowing it to adapt and improve over time [47, 25, 52]. In this approach, PINNs optimized with the NSGA-III algorithm can be incrementally updated with new observational data. This continual learning process allows the model to refine its predictions, improve accuracy, and maintain robustness when additional data becomes available. Moreover, it facilitates the dynamic update of systems and uncertainties, making it particularly valuable for real-time prediction and long time simulation of complex physical systems [18, 54, 61].

### v2

@@ 3.2.3 Inverse Problem
Table 10 shows the estimated $\alpha$ values of the TFMDWEs and their $L^{1}$ error in the inverse problem setting for (1). ADAM-PINN, (2). NSGA-III-PINN, and (3). iPINNER with different noise level of observational data. It is shown that in all cases, iPINNER is more accurate in estimating the missing parameters in TFMDWEs than ADAM-PINN and NSGA-III-PINN. Similar as in the forward problem setting, the NSGA-III-PINN shows more accurate estimations than ADAM-PINN, as the NSGA-III algorithm more effectively balances multiple objectives within the loss function, thereby partially mitigating the adverse effects of noisy observations even in the missing parameters. This is also reflected in the figures 1113 as well as Figure 14 which shows at different time instances (top: $t=0.5$, bottom: $t=1$), the true solution (black) and inverse problem solutions obtained using ADAM-PINN (red), NSGA-III-PINN (blue), and iPINNER (green) for the time-fractional mixed diffusion-wave equations (TFMDWEs) with different noise levels. The results show that iPINNER (green) outperforms both ADAM-PINN (red) and NSGA-III-PINN (blue) in terms of estimating parameters and PDE solutions.
Figure 11: Comparison of (a) the true solution and (b) the iPINNER solution as well as the $L_{1}$ errors of (c) the ADAM-PINN and (d) the iPINNER for the time-fractional mixed diffusion-wave equations (TFMDWEs) in inverse problem. Observation noise level is $20\%$.
Figure 12: Comparison of (a) the true solution and (b) the iPINNER solution as well as the $L_{1}$ errors of (c) the ADAM-PINN and (d) the iPINNER for the time-fractional mixed diffusion-wave equations (TFMDWEs) in inverse problem. Observation noise level is $50\%$.
Figure 13: Comparison of (a) the true solution and (b) the iPINNER solution as well as the $L_{1}$ errors of (c) the ADAM-PINN and (d) the iPINNER for the time-fractional mixed diffusion-wave equations (TFMDWEs) in inverse problem. Observation noise level is $80\%$.
@@
Figure 14: Comparison at different time instances (top: $t=0.5$, bottom: $t=0.7$) between the true solution (black) and inverse problem solutions obtained using ADAM-PINN (red), NSGA-III-PINN (blue), and iPINNER (green) for the time-fractional mixed diffusion-wave equations (TFMDWEs) under varying noise levels: Panel (a) — $20\%$ noise, Panel (b) — $50\%$ noise, and Panel (c) — $80\%$ noise.
@@
iPINNER 0.5713 0.0713
Table 10: Comparison of PINN with Adam optimizer (ADAM-PINN), PINN with NSGA-III optimizer (NSGA-III-PINN) and the proposed integrated method (iPINNER) for estimating the fractional term $\alpha$ in the TFMDWEs. For reference, the true value of the fractional term is $\alpha=0.5$.
3.3 Two-Dimensional Heat Equation
We next consider the classical two-dimensional heat equation on the unit square with homogeneous Dirichlet boundaries:
$\partial_{t}u(x,y,t)-\kappa\,\Delta u(x,y,t)\;=\;f(x,y,t),\qquad(x,y)\in\Omega\equiv[0,1]^{2},\;t\in[0,1],$ (25)
$u|_{\partial\Omega}=0,$ (26)
$u(x,y,0)=u_{0}(x,y).$ (27)
To enable quantitative comparison, we consider the analytical solution
$u(x,y,t)=\sin(\pi x)\sin(\pi y)\,\mathrm{e}^{-2\pi^{2}\kappa t},$ (28)
which corresponds to non forcing case, i.e., $f\equiv 0$ and the initial condition $u_{0}(x,y)=\sin(\pi x)\sin(\pi y)$.
Neural network architecture
In order to make a fair comparison, with different optimizers, the PINN employs a consistent deep neural network architecture comprising 6 layers with 64 neurons in each layer. During the training phase, models corresponding to each optimizer were collected until the training loss converged below a prescribed threshold $\epsilon$. Specifically, the ADAM-PINN requires $30000$ epochs to achieve convergence. The proposed iPINNER requires of $5$ generations with $5000$ epochs to ensure the convergence of training loss.
Training and Testing Data
The training data for the 2D Heat equation consist of three main components: Initial Condition (IC) points, Boundary Condition (BC) points, and Collocation Points. The IC points are sampled in space at the initial time $t=0$, while the BC points are sampled in time along the spatial boundaries $\partial\Omega$. The collocation points are randomly selected from the interior of the spatiotemporal domain, with 100 points used in this study. The testing data are sampled over the domain $\Omega\times[0,T]$ with $\Omega=[0,1]$. The spatial mesh size is of $h=\Delta x=\Delta y=1/10$ and a temporal step size is $\Delta t=0.01$.
Error criteria
We evaluate models using the $L_{2}$ relative error between the benchmark solutions and the predictions of different models.
In the forward problem setting, we samples the observational data at $N_{\text{obs}}$ scattered space–time locations, which are then added with zero-mean Gaussian noise with standard deviation equal to a prescribed fraction $\eta\in\{20\%,50\%\}$ of $\mathrm{std}(u)$ evaluated at those points (consistent with Sections 3.13.2).
Figure 15: iPINNER prediction of 2D Heat equation with 20% noise data.
Figure 16: iPINNER prediction of 2D Heat equation with 50% noise data.
Figures 1516 show the forward problem solutions using iPINNER (left) and the corresponding true solutions (middle) of the 2D heat equation at two different time instances (top: $t=0.5$, bottom: $t=1$) under different noise levels. The right panels show the corresponding $L^{1}$ errors between the predicted and true solutions. iPINNER shows higher accuracy than ADAM-PINN, particularly under higher noise levels. With 50% data noise, the iPINNER has 0.0180 $L_{2}$ relative error whereas ADAM-PINN has 0.0285 $L_{2}$ relative error shown in Table 11
Table 11: Wall time of training PINN and iPINNER with 50% noise data.
Method Wall Time (s) $L_{2}$ relative error
Standard PINN 256 0.0285
iPINNER 430 0.0180
@@ 4 Conclusion and Future Work
In this paper, we introduce a novel iPINNER framework that integrates the physics informed neural network (PINN) with ensemble Kalman filter (EnKF) with the NSGA-III multi-objective optimizer. This framework can be used to address both forward problem and inverse problem with only limited noisy data in the context of partial differential equations (PDE). The iPINNER framework has several advantages: (i) The framework utilizes the multi-objective optimizer NSGA-III to find an optimal cluster of PINNs, and (ii) The uncertainty within this cluster, inherent from the NSGA-III optimization, can be integrated with observational noisy data using the ensemble Kalman filter. This process can then be iteratively applied to update the data loss during PINN training. (iii) The framework can be applied in both the forward and inverse problems in solving PDEs.
Building on the iPINNER framework’s effectiveness in solving forward and inverse problems with noisy data, several future directions necessitate further investigation. First, in the inverse problem setting, the unknown parameters are treated as trainable variables within the neural networks and are implicitly embedded in the PDE residual loss of the PINN. To quantify the uncertainty of the inferred parameters, these parameters together with the independent variables can be used as inputs to the neural network, allowing it to be trained across a wide range of physical parameters [harlim2021machine, gottwald2021combining]. Second, a Bayesian PINN framework for both inverse and forward models, as proposed in [yang2021b], can be integrated with the current framework. This integration could potentially improve the performance of the model in the presence of highly noisy data, as the Bayesian approach naturally quantifies the uncertainties arising from scattered noisy data. Third, the iPINNER framework can be extended to a continual learning framework, allowing it to adapt and improve over time [purvine2017comparative, howard2024multifidelity, roy2024exact]. In this approach, PINNs optimized with the NSGA-III algorithm can be incrementally updated with new observational data. This continual learning process allows the model to refine its predictions, improve accuracy, and maintain robustness when additional data becomes available. Moreover, it facilitates the dynamic update of systems and uncertainties, making it particularly valuable for real-time prediction and long time simulation of complex physical systems [ghunaim2023real, shaheen2022continual, yang2012adaptive].
@@ Acknowledgment
Appendix A Mathematical Formulation of Ensemble Kalman Filter (EnKF)
Consistent with the notation in Section 2.3, let the forecast ensemble at time $t$ be denoted as follows
$X_{t|t-1}=\bigl[x_{t|t-1}^{(1)},,x_{t|t-1}^{(2)},,\dots,,x_{t|t-1}^{(N)}\bigr]\in\mathbb{R}^{d\times N},$ (29)
where $N$ is the ensemble size and $d$ denotes the dimension of the state variable. Each column $x_{t|t-1}^{(i)}$ represents one realization of the state vector in forecast models. Then, the prior mean $\bar{x}_{t|t-1}$ and covariance $R_{t|t-1}$ (model forecasts) are then estimated by the following [evensen2003ensemble]
$\bar{x}_{t|t-1}$ $=\frac{1}{N}\sum{i=1}^{N}x_{t|t-1}^{(i)},$ (30)
$R_{t|t-1}$ $=\frac{1}{N-1}\sum_{i=1}^{N}\bigl(x_{t|t-1}^{(i)}-\bar{x}_{t|t-1}\bigr)\bigl(x_{t|t-1}^{(i)}-\bar{x}_{t|t-1}\bigr)^{\top},$ (31)
which provide Monte Carlo approximations to the prior mean and covariance. Then, given an observation $y_{t}$, the ensemble members can be updated through Bayesian formula and yields the following:
$x_{t}^{(i)}=x_{t|t-1}^{(i)}+K_{t}\bigl(y_{t}^{(i)}-Hx_{t|t-1}^{(i)}\bigr),\quad i=1,\dots,N,$ (32)
where $y_{t}^{(i)}=y_{t}+\epsilon_{t}^{(i)}$ are observations with observation error $\epsilon_{t}^{(i)}\sim\mathcal{N}(0,R)$. The ensemble Kalman gain can be derived as follows:
$K_{t}=P_{t|t-1}H^{\top}\bigl(HP_{t|t-1}H^{\top}+R\bigr)^{-1}.$ (33)